OpenStudy (candycane):

what is the simpler form of each product: (8m+1)^2 the answer id 64m^2+16m+1, now I get where the 64m^2 and the 1 came from, Im confused about how 16m came to be part of the equation, can someone please explain?

OpenStudy (anonymous):

(a+b)^2 = (a+b)*(a+b) Applying distributive rule (a+b)^2 = (a+b)*a + (a+b)*b (a+b)^2 = a*a+b*a + a*b+b*b Note that a*a=a^2 and a*b=b*a (a+b)^2 = a^2+ 2*(a*b) + b^2 So: (8m+1)^2 = (8m)^2+ 2*((8m)*1) + (1)^2 (8m+1)^2 = (8^2)*(m^2)+ 2*8*m*1 + 1 (8m+1)^2 = (64)*(m^2)+ 16*m + 1 (8m+1)^2 = 64m^2+ 16m + 1 Hope that helps!

OpenStudy (candycane):

Yes that helps, Thanks :)